Please solve this using the substitution method.
x = 3y - 5
6x - 18y = -30

Answer by checkley77(12844)

x = 3y - 5
6x - 18y = -30
6(3y-5)-18y=-30
18y-30-18y=-30
0y=-30+30
0y=0
There is no unique solution to these 2 equations for the x & y terms.
The 2 equations are identical.
Should you plot these 2 equations, you'd find that they form a single line which has an infinite number of solutions for x & y.